# skipabove=5pt, innertopmargin=0pt

### v1

@@ Abstract
As large language models (LLMs) are becoming more capable and widespread, the study of their failure cases is becoming increasingly important. Recent advances in standardizing, measuring, and scaling test-time compute suggest new methodologies for optimizing models to achieve high performance on hard tasks. In this paper, we apply these advances to the task of “model jailbreaking”: eliciting harmful responses from aligned LLMs. We develop an adversarial reasoning approach to automatic jailbreaking via test-time computation that achieves SOTA attack success rates (ASR) against many aligned LLMs, even the ones that aim to trade inference-time compute for adversarial robustness. Our approach introduces a new paradigm in understanding LLM vulnerabilities, laying the foundation for the development of more robust and trustworthy AI systems. Code is available at https://github.com/Helloworld10011/Adversarial-Reasoning.
@@ 1 Introduction
Large language models (LLMs) are increasingly deployed with various safety techniques to ensure alignment with human values. Common strategies include RLHF [CLB+23, OWJ+22], DPO [RSM+24], and the usage of dedicated guardrail models [IUC+23, RDS+23]. In nominal use cases, alignment methods typically refuse to generate objectionable content, but adversarially designed prompts can bypass these guardrails. A challenge known as jailbreaking consists of finding prompts that circumvent safety measures and elicit undesirable behaviors.
Current jailbreaking methods fall into two categories: token-space and prompt-space attacks. Token-space attacks [SRI+20, WJK+23, ZWC+23, HBC+24, ACF24] focus on token-level modifications of the input to minimize some loss value, often using gradient-based heuristics [ZWC+23] or random searches [ACF24]. Such methods view jailbreaking as an optimization problem over sequences of tokens and use the loss information to inform their navigation through the optimization landscape. As a consequence, token-level methods produce semantically meaningless input prompts that can be mitigated by perplexity-based or smoothing-based filters [AK23, RWHP24].
In contrast, prompt-space attacks generate semantically coherent adversarial prompts using techniques like multi-round LLM interactions or fine-tuning for crafted outputs [CRD+24, GZH+23, LXCX24, MZK+24, ZLZ+24, SRL+24, LLS+24b]. A notable subset deploys chain-of-thought reasoning [WWS+23, NAGA+21] to guide the interaction with the target LLM and better navigate the prompt search space [CRD+24, MZK+24]. These methods are designed to exploit binary feedback from the target LLM of the form “has the current prompt jailbroken the target LLM or not?”. Binary feedback effectively quantifies attack success on black-box models, but it provides limited information for the intermediate stages. This sparsity hampers search effectiveness in the prompt space to exploit the safety vulnerabilities of the target LLM, particularly against adversarially trained models [ZPW+24, SEG+24, XSG+24]. An extensive search through the prompt space requires a more granular signal—e.g., loss values—to efficiently traverse the prompt landscape and identify the target LLM weaknesses.
@@
  • In this paper, we formulate jailbreaking as a reasoning problem. We then apply insights from the reasoning field and lessons from existing token-space and prompt-space methods to create a strong, adaptive, gradient-free, and transferable jailbreaking algorithm.
  • Experimentally, we show that our method achieves state-of-the-art attack success rates among prompt-space attacks and even outperforms token-space methods for many target LLMs, particularly those that have been adversarially trained (see Table 1). Notably, our method enhances the jailbreaking performance significantly when a weak LLM is used as the attacker (Table 2), reflecting the benefits of optimizing test-time computation. We further introduce a multi-shot transfer attack scenario that outperforms existing methods and achieves 56% success rate on OpenAI o1-preview (Table 3) and 100% on Deepseek. Finally, in our ablation studies, we show that our method (i) effectively reduces the loss (Figure 3) , and quantitatively demonstrate the key role of the feedback in this regard (Figure 4); (ii) we also show that our method benefits from deeper reasoning, i.e., it continues to discover new jailbreaks with more iterations. (Figure 5).
@@ 2 Related Work
Token-space Jailbreaking. Token-space attacks [SRI+20, WJK+23, ZWC+23, HBC+24, ACF24] modify the input at the token level to decrease some loss value. For example, the GCG algorithm [ZWC+23], one of the first transferrable token-level attacks to achieve significant success rates on aligned models, uses the gradient of the loss to guide the greedy search. Subsequent work has refined this approach to obtain lower computational cost and improved effectiveness [LS24, JPD+24], including token-level modifications by other heuristics and without a gradient [HBC+24] and random searches over cleverly chosen initial prompts [ACF24]. We adopt the use of a loss function from these methods as a signal to inform how to navigate the prompt-space for better jailbreaks while remaining gradient-free.
Prompt-space Jailbreaking. These methods often rely on a "red-teaming" LLM to generate adversarial prompts [PHS+22, WHS23, SSS+24, CRD+24, LXCX24, MZK+24, ZLZ+24, SRL+24, LLS+24b]. Methods such as PAIR [CRD+24] deploy a separate LLM, called the "attacker", which uses a crafted system prompt to interact with the target LLM over multiple rounds and generate semantic jailbreaks; they operate in a black-box manner, requiring only the target’s outputs, and are highly transferrable [CRD+24]. Some other methods fine-tune a model to generate the attacking prompts [PHS+22, GZH+23, ZLZ+24, PZG+24, BCW+24], though this demands substantial computational resources. Rather than fine-tuning, we rely on increased test-time computation [SLXK24], while others start from expert-crafted prompts (e.g., DAN [Cha24]) and refine them via genetic algorithms [LXCX24, SRL+24, LLS24a]. Like these methods, our approach generates semantically meaningful jailbreaks by using another LLM as the attacker, however, our approach is significantly different from the prior work as we develop reasoning modules based on the loss values to better navigate the prompt space.
@@
Comparison with PAIR and TAP. The closest methods to our framework are PAIR [CRD+24] and TAP-T [MZK+24]. Our verifier-driven method outperforms PAIR and TAP, whose prompt-based CoT remains static and does not leverage a loss. That said, while TAP-T creates a tree of attacks based on the attacker’s CoT, it prunes only prompts that do not request the same content as the original intent, and does not utilize any reasoning methodologies.
@@
Chain-of-Thought (CoT). CoT prompting [WWS+23] and scratch-padding [NAGA+21] demonstrate how prompting the model to produce a step-by-step solution improves the LLM’s performance. Recent work [ZLC+24, WZW+24, XSG+25] suggests constructing the CoT through several modules rather than relying only on the language model’s CoT capabilities. Notably, [XSG+25] explicitly constructs the CoT to ensure that it follows a particular path. Likewise, we use three modules for explicitly constructing the reasoning steps, aiming to reduce the loss function with each step.
@@
Reasoning. Recent advances in reasoning have enhanced LLMs’ capabilities in solving complex problems by scaling test-time computation mechanisms [HBK+21, RPBN+23, AVL+24, SWZ+24, RHS+23, Ope24b]. “Best-of-N” sampling [CKB+21, YGW24], which runs N parallel streams and verifies the final answers through outcome-based reward models (ORMs), is a straightforward test-time scaling approach. However, it might fail to uncover solutions that require incremental improvements, limiting its effectiveness compared to other test-time methods. To address this limitation, recent work utilizes process-based reward models (PRMs) [UKK+22, LKB+23, WLS+24] to optimally scale the test-time computation, thereby improving the reasoning performance [SLXK24, XGZ+24, GLG+24, Ope24a]. PRMs provide step-by-step verification that facilitates a look-ahead signal at each step, often necessary for a searching algorithm [XGZ+24]. Similarly, the use of a continuous loss function as a step-wise verifier allows us to run a tree search. This framework fits well into the “Proposer and Verifier” perspective [SLXK24] of test-time computation, where the proposer proposes a distribution of solutions and a verifier assigns rewards to the proposed distributions. Robust verifier models are essential for accurate guidance [ZKL+24, ZHB+24, SVK24], but they require intermediate annotations from human annotators or heuristics [UKK+22, LKB+23, WLS+24]. In our work, we use the loss values from a surrogate LLM as a verifier, eliminating the need for training a verifier model.
@@
Reasoning vs safety. The reasoning framework for exploiting test-time compute can also be used to improve alignment. OpenAI uses “deliberative alignment” [GJW+24] to incorporate human-generated and adversarial data to improve the alignment of the o1 model family [Ope24b, Ope24a]. These models consistently outperform traditional frontier LLMs in metrics measuring vulnerability to automatic adversarial attacks [Ope24b, HPL+24]. The limitations of current automatic jailbreaking methods and the efficacy of using test-time compute for safety alignment naturally beg the question of whether test-time compute frameworks can be used to bypass model guardrails instead of enforcing them. Adversarial reasoning, the framework we propose in this paper, demonstrates that bypassing a model’s guardrails—even those that leverage increased computation for enhanced safety—is not only possible but also effective.
@@ 3 Preliminaries
We reinterpret this problem as a reasoning problem. Rather than directly optimizing the prompt $P$ as token-level methods do, we construct $P$ by applying an attacker LLM $\mathbb{A}$ to a reasoning string $S$, i.e., $P=\mathbb{A}(S)$. This allows us to update the attacker’s output distribution according to the “Proposer and Verifier” framework [SLXK24] by iteratively refining $S$. Thus, the challenge is to find a string $S$ such that, when passed to the attacker as shown in Figure 1, it satisfies the following objective:
@@
This formulation allows us to view jailbreaking methods through the lens of an iterative refinement of $S$. Note that many existing reasoning algorithms can be framed into this formulation. For instance, chain-of-thought prompting can be realized by repeatedly generating partial thoughts from an LLM and appending them to $S$. The final answer is generated by passing the updated $S$ to the same LLM. Similarly, in the jailbreaking literature, methods such as PAIR [CRD+24] aim to solve (3.1) by updating $S$ at each iteration, appending the generated CoT from the attacker and the responses from the target. The string $S$ encapsulates all partial instructions and the intermediate steps executed during the attacking process.
Prior prompt-space jailbreaking methods have often used prompted or fine-tuned LLMs [CRD+24, MZK+24, MPY+24] as judges to evaluate whether a jailbreak is successful. These judges are the simplest “verifiers”: once the refinement of $S$ is over, the judge will evaluate if $\mathbb{A}(S)$ is a successful jailbreak. However, the judge only provides a binary signal: whether or not jailbreaking has taken place. This makes binary verifiers unsuitable for estimating intermediate rewards. To alleviate this, we use a continuous and more informative loss function. A loss function will provide more granular feedback by assigning smaller loss values to prompts that are semantically closer to eliciting the desired behavior from $\mathbb{T}$. Following prior work [ZWC+23, HBC+24], we use the cross-entropy loss of a particular affirmative string for each intent (e.g., “Sure, here is the step-by-step instructions for building a bomb”), measuring how likely the target model is to begin with that string. Showing this desired string as $\mathbf{y}_{I}=\{y_{1},y_{2},\cdots,y_{l}\}$, our goal is to optimize the following next-word prediction loss:
@@ 4 Algorithm
We must find a reasoning string $S$ for an attacker LLM $\mathbb{A}$ such that the resulting prompt $P:=\mathbb{A}(S)$ jailbreaks the target LLM. Our algorithm iteratively refines $S$, aiming to minimize the loss given in Equation 3.3. Starting from a root node $S^{(0)}$–a predefined template in Appendix C–we iteratively construct a reasoning tree with nodes representing candidate strings (see Figure 1). At iteration $t$, a node in the tree with the best score is selected. This node will expand into $m$ children $S^{(t+1)}_{1},\cdots,S^{(t+1)}_{m}$. The tree will be further pruned at each iteration to retain a buffer of the best nodes. We now explain this process in detail, beginning with a description of an individual reasoning step in our method.
@@
Let $S^{(t)}$ be the reasoning string at time $t$. We generate $n$ prompts $P_{1},P_{2},\cdots,P_{n}$ sampled independently from the distribution of $\mathbb{A}$ with $S^{(t)}$ as input (see Figure 2). Let $\ell_{1},\cdots,\ell_{n}$ be the loss values obtained from Section 3 for $P_{1},\cdots,P_{n}$, respectively. For simplicity, we assume that the prompts are ordered in a way that $\ell_{1}\leq\ell_{2}\leq\dots\leq\ell_{n}$, i.e., prompt $P_{1}$ incurs the lowest loss and $P_{n}$ the highest loss on the target LLM. We use Feedback LLM $\mathbb{F}$ with a crafted system prompt (given in Appendix C) that takes the ordered set of prompts as input and generates a feedback string $F$ as a textual analysis, semantically explaining why $P_{1}$ is a better attacking prompt than $P_{2}$, why $P_{2}$ is better than $P_{3}$, etc., and identifies any apparent patterns across the prompts:
@@
Once the feedback is obtained, the reasoning string must be refined and updated into its children. One way to do this is to append the feedback to the reasoning string at each time. This quickly becomes intractable—not only does the string length grow with each iteration, but also the set of different feedbacks can contradict each other. Instead, we deploy a Refiner LLM $\mathbb{R}$ inspired by the idea of “textual gradients” introduced in [YBB+24]. Taking $S^{(t)}$ and the feedback string $F$ as its arguments, $\mathbb{R}$ generates a new reasoning string $S^{(t+1)}$ that refines $S^{(t)}$ based on the feedback:
@@
Replicating the above process (Feedback+Refine) $m$ times in parallel, we produce $m$ new reasoning strings $\{S^{(t+1)}_{1},\cdots,S^{(t+1)}_{m}\}$ as the children of $S^{(t)}$. Figure 2 shows a single iteration of our method and illustrates how the updated reasoning string incorporates the key components of the feedback. Note that rather than relying on the attacker’s CoT process—which lacks any information about the loss function—we explicitly engineer the reasoning steps aiming to decrease the loss function. This setup parallels recent efforts that align a model’s intermediate steps with predefined strategies [WZW+24, XSG+25].
@@
We now describe the node selection and pruning process based on the verifier’s score. Inspired by the Go-with-the-Winners algorithm [AV94] and Greedy Coordinate Query [HBC+24], our method maintains a priority buffer of reasoning strings with the highest scores according to Equation 4.3, pruning those with lower scores (i.e., strings that generate attacking prompts with high loss). The buffer is denoted as $L$ in Algorithm 1 with a size of $B$. At each iteration, the highest-scoring node in the buffer is selected for expansion, and its $m$ children are added to the buffer. The buffer is then pruned to retain only the top $B$ candidates. This pruning strategy enables backtracking when none of the children exceed the scores of existing buffer candidates. Algorithm 1 outlines these steps. In line 4, the reasoning string is initialized with a template and iteratively updated to generate new attacking prompts. This process repeats for $T$ iterations.

### v2

# Adversarial Reasoning at Jailbreaking Time

@@ Abstract
As large language models (LLMs) are becoming more capable and widespread, the study of their failure cases is becoming increasingly important. Recent advances in standardizing, measuring, and scaling test-time compute suggest new methodologies for optimizing models to achieve high performance on hard tasks. In this paper, we apply these advances to the task of “model jailbreaking”: eliciting harmful responses from aligned LLMs. We develop an adversarial reasoning approach to automatic jailbreaking that leverages a loss signal to guide the test-time compute, achieving SOTA attack success rates against many aligned LLMs, even those that aim to trade inference-time compute for adversarial robustness. Our approach introduces a new paradigm in understanding LLM vulnerabilities, laying the foundation for the development of more robust and trustworthy AI systems. Code is available at Github.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) are increasingly deployed with various safety techniques to ensure alignment with human values. Common strategies include RLHF (Christiano et al., 2017; Ouyang et al., 2022), DPO (Rafailov et al., 2024), and the usage of dedicated guardrail models (Inan et al., 2023; Rebedea et al., 2023). In nominal use cases, alignment methods typically refuse to generate objectionable content, but adversarially designed prompts can bypass these guardrails. A challenge known as jailbreaking consists of finding prompts that circumvent safety measures and elicit undesirable behaviors.
Current jailbreaking methods fall into two categories: token-space and semantic-space attacks. Token-space attacks (Shin et al., 2020; Wen et al., 2023; Zou et al., 2023; Hayase et al., 2024; Andriushchenko et al., 2024a) focus on token-level modifications of the input to minimize some loss value, often using gradient-based heuristics (Zou et al., 2023) or random searches (Andriushchenko et al., 2024a). Such methods view jailbreaking as an optimization problem over sequences of tokens and use the loss information to inform their navigation through the optimization landscape. As a consequence, token-level methods produce semantically meaningless input prompts that can be mitigated by perplexity-based or smoothing-based filters (Alon & Kamfonas, 2023; Robey et al., 2024b).
In contrast, semantic-space attacks generate semantically coherent adversarial prompts using techniques like multi-round LLM interactions or fine-tuning for crafted outputs (Chao et al., 2024b; Ge et al., 2023; Liu et al., 2024b; Mehrotra et al., 2024; Zeng et al., 2024; Samvelyan et al., 2024; Liu et al., 2024a). A notable subset deploys chain-of-thought reasoning (Wei et al., 2023b; Nye et al., 2021) to guide the interaction with the target LLM and better navigate the prompt search space (Chao et al., 2024b; Mehrotra et al., 2024). These methods are designed to exploit binary feedback from the target LLM of the form “has the current prompt jailbroken the target LLM or not?”. Binary feedback effectively quantifies attack success on black-box models, but it provides limited information for the intermediate stages. This sparsity hampers search effectiveness in the prompt space to exploit the safety vulnerabilities of the target LLM, particularly against adversarially trained models (Zou et al., 2024; Sheshadri et al., 2024; Xhonneux et al., 2024). An extensive search through the prompt space requires a more granular signal—e.g., loss values—to efficiently traverse the prompt landscape and identify the target LLM weaknesses.
@@
  • In this paper, we formulate jailbreaking as a reasoning problem. We then apply insights from the reasoning field and lessons from existing token-space and semantic methods to create a strong, adaptive, gradient-free, and transferable jailbreaking algorithm.
  • Experimentally, we show that our method achieves state-of-the-art success rates among semantic-space attacks and outperforms token-space methods for many target LLMs, particularly those that have been adversarially trained (see Table 1). Notably, our method enhances the jailbreaking performance significantly when a weak LLM is used as the attacker (Table 3), reflecting the benefits of optimizing test-time computation. We further introduce a multi-shot transfer scenario—as the method relies on the target’s logit vectors—that outperforms existing methods and achieves 56% success rate on OpenAI o1-preview (Table 4) and 100% on Deepseek. Finally, in our ablation studies, we show that (i) our method effectively reduces the loss (Figure 3) and quantitatively demonstrate the key role of the feedback (Figure 4); (ii) a frontier reasoning model such as Deepseek-R1 (DeepSeek-AI, 2025) as the attacker does not gain additional success when working heuristically and not supervised; and (iii) our method benefits from deeper reasoning, i.e., it continues to discover new jailbreaks with more iterations. (Figure 5).
@@ 2 Related Work
Comparison with PAIR and TAP. The closest methods to our framework are PAIR (Chao et al., 2024b) and TAP-T (Mehrotra et al., 2024). Our verifier-driven method outperforms PAIR and TAP that rely on LLM’s inherent CoT reasoning and does not leverage a loss. That said, while TAP-T creates a tree of attacks based on the attacker’s CoT, it prunes only prompts that do not request the same content as the original intent, and does not utilize any reasoning methodologies.
@@
Chain-of-Thought (CoT). CoT prompting (Wei et al., 2023b) and scratch-padding (Nye et al., 2021) demonstrate how prompting the model to produce a step-by-step solution improves the LLM’s performance. Recent work (Zheng et al., 2024; Wang et al., 2024b; Xiang et al., 2025) suggests constructing the CoT through several modules rather than relying only on the language model’s CoT capabilities. Notably, (Xiang et al., 2025) explicitly constructs the CoT to ensure that it follows a particular path. Likewise, we use three modules for explicitly constructing the reasoning steps, aiming to reduce the loss function with each step.
@@
Reasoning. Recent advances in reasoning have enhanced LLMs’ capabilities in solving complex problems by scaling test-time computation mechanisms (Hendrycks et al., 2021; Romera-Paredes et al., 2023; Ahn et al., 2024; Shao et al., 2024; Rein et al., 2023; OpenAI, 2024b). “Best-of-N” sampling (Cobbe et al., 2021; Yu et al., 2024), which runs N parallel streams and verifies the final answers through outcome-based reward models (ORMs), is a straightforward test-time scaling approach. However, it might fail to uncover solutions that require incremental improvements, limiting its effectiveness compared to other test-time methods. To address this limitation, recent work utilizes process-based reward models (PRMs) (Uesato et al., 2022; Lightman et al., 2023; Wang et al., 2024a) to optimally scale the test-time computation, thereby improving the reasoning performance (Snell et al., 2024; Xie et al., 2024; Gandhi et al., 2024; OpenAI, 2024a). PRMs provide step-by-step verification that facilitates a look-ahead signal at each step, often necessary for a searching algorithm (Xie et al., 2024). Similarly, the use of a continuous loss function as a step-wise verifier allows us to run a tree search. This framework fits well into the “Proposer and Verifier” perspective (Snell et al., 2024) of test-time computation, where the proposer proposes a distribution of solutions and a verifier assigns rewards to the proposed distributions. Robust verifier models are essential for accurate guidance (Zhang et al., 2024a, b; Stechly et al., 2024), but they require intermediate annotations from human annotators or heuristics (Uesato et al., 2022; Lightman et al., 2023; Wang et al., 2024a). In our work, we use the loss values from a surrogate LLM as a verifier, eliminating the need for training a verifier model.
@@
Reasoning vs safety. The reasoning framework for exploiting test-time compute can also be used to improve alignment. OpenAI uses “deliberative alignment” (Guan et al., 2025) to incorporate human-generated and adversarial data to improve the alignment of the o1 model family (OpenAI, 2024a, b). These models consistently outperform traditional frontier LLMs in metrics measuring vulnerability to automatic adversarial attacks (OpenAI, 2024b; Hughes et al., 2024). The limitations of current automatic jailbreaking methods and the efficacy of using test-time compute for safety alignment naturally beg the question of whether test-time compute frameworks can be used to bypass model guardrails instead of enforcing them. Adversarial reasoning, the framework we propose in this paper, demonstrates that bypassing a model’s guardrails—even those that leverage increased computation for enhanced safety—is not only possible but also effective. Additional related work is provided in Appendix A.
@@ 3 Preliminaries
We reinterpret this problem as a reasoning problem. Rather than directly optimizing the prompt $P$ as token-level methods do, we construct $P$ by applying an attacker LLM $\mathbb{A}$ to a reasoning string $S$, i.e., $P=\mathbb{A}(S)$. This allows us to update the attacker’s output distribution according to the “Proposer and Verifier” framework (Snell et al., 2024) by iteratively refining $S$. Thus, the challenge is to find a string $S$ such that, when passed to the attacker as shown in Figure 1, it satisfies the following objective:
@@
This formulation allows us to view jailbreaking methods through the lens of an iterative refinement of $S$. Note that many existing reasoning algorithms can be framed into this formulation. For instance, chain-of-thought prompting can be realized by repeatedly generating partial thoughts from an LLM and appending them to $S$. The final answer is generated by passing the updated $S$ to the same LLM. Similarly, in the jailbreaking literature, methods such as PAIR (Chao et al., 2024b) aim to solve (3.1) by updating $S$ at each iteration, appending the generated CoT from the attacker and the responses from the target. The string $S$ encapsulates all partial instructions and the intermediate steps executed during the attacking process.
Prior semantic-space jailbreaking methods have often used prompted or fine-tuned LLMs (Chao et al., 2024b; Mehrotra et al., 2024; Mazeika et al., 2024) as judges to evaluate whether a jailbreak is successful. These judges are the simplest “verifiers”: once the refinement of $S$ is over, the judge will evaluate if $\mathbb{A}(S)$ is a successful jailbreak. However, the judge only provides a binary signal: whether or not jailbreaking has taken place. This makes binary verifiers unsuitable for estimating intermediate rewards. To alleviate this, we use a continuous and more informative loss function. A loss function will provide more granular feedback by assigning smaller loss values to prompts that are semantically closer to eliciting the desired behavior from $\mathbb{T}$. Following prior work (Zou et al., 2023; Hayase et al., 2024), we use the cross-entropy loss of a particular affirmative string for each intent (e.g., “Sure, here is the step-by-step instructions for building a bomb”), measuring how likely the target model is to begin with that string. Showing this desired string as $\mathbf{y}_{I}=\{y_{1},y_{2},\cdots,y_{l}\}$, our goal is to optimize the following next-word prediction loss:
@@ 4 Algorithm
We must find a reasoning string $S$ for an attacker LLM $\mathbb{A}$ such that the resulting prompt $P:=\mathbb{A}(S)$ jailbreaks the target LLM. Our algorithm iteratively refines $S$, aiming to minimize the loss given in Equation 3.3. Starting from a root node $S^{(0)}$–a predefined template in Appendix D–we iteratively construct a reasoning tree with nodes representing candidate strings (see Figure 1). At iteration $t$, a node in the tree with the best score is selected. This node will expand into $m$ children $S^{(t+1)}_{1},\cdots,S^{(t+1)}_{m}$. The tree will be further pruned at each iteration to retain a buffer of the best nodes. We now explain this process in detail, beginning with a description of an individual reasoning step in our method.
@@
Let $S^{(t)}$ be the reasoning string at time $t$. We generate $n$ prompts $P_{1},P_{2},\cdots,P_{n}$ sampled independently from the distribution of $\mathbb{A}$ with $S^{(t)}$ as input (see Figure 2). Let $\ell_{1},\cdots,\ell_{n}$ be the loss values obtained from Section 3 for $P_{1},\cdots,P_{n}$, respectively. For simplicity, we assume that the prompts are ordered in a way that $\ell_{1}\leq\ell_{2}\leq\dots\leq\ell_{n}$, i.e., prompt $P_{1}$ incurs the lowest loss and $P_{n}$ the highest loss on the target LLM. We use Feedback LLM $\mathbb{F}$ with a crafted system prompt (given in Appendix D) that takes the ordered set of prompts as input and generates a feedback string $F$ as a textual analysis, semantically explaining why $P_{1}$ is a better attacking prompt than $P_{2}$, why $P_{2}$ is better than $P_{3}$, etc., and identifies any apparent patterns across the prompts:
@@
Once the feedback is obtained, the reasoning string must be refined and updated into its children. One way to do this is to append the feedback to the reasoning string at each time. This quickly becomes intractable—not only does the string length grow with each iteration, but also the set of different feedbacks can contradict each other. Instead, we deploy a Refiner LLM $\mathbb{R}$ inspired by the idea of “textual gradients” introduced in (Yuksekgonul et al., 2024). Taking $S^{(t)}$ and the feedback string $F$ as its arguments, $\mathbb{R}$ generates a new reasoning string $S^{(t+1)}$ that refines $S^{(t)}$ based on the feedback:
@@
Replicating the above process (Feedback+Refine) $m$ times in parallel, we have $m$ new reasoning strings $\{S^{(t+1)}_{1},\cdots,S^{(t+1)}_{m}\}$ as the children of $S^{(t)}$. Figure 2 shows a single iteration of our method and illustrates how the updated reasoning string incorporates the key components of the feedback. Note that rather than relying on the attacker’s CoT process—which lacks any information about the loss function—we explicitly engineer the reasoning steps aiming to decrease the loss function. This setup parallels recent efforts that align a model’s intermediate steps with predefined strategies (Wang et al., 2024b; Xiang et al., 2025).
@@
We now describe the node selection and pruning process based on the verifier’s score. Inspired by the Go-with-the-Winners algorithm (Aldous & Vazirani, 1994) and Greedy Coordinate Query (Hayase et al., 2024), our method maintains a priority buffer of reasoning strings with the highest scores according to Equation 4.3, pruning those with lower scores (i.e., strings that generate attacking prompts with high loss). The buffer is denoted as $L$ in Algorithm 1 with a size of $B$. At each iteration, the highest-scoring node in the buffer is selected for expansion, and its $m$ children are added to the buffer. The buffer is then pruned to retain only the top $B$ candidates. This pruning strategy enables backtracking when none of the children exceed the scores of existing buffer candidates. Algorithm 1 outlines these steps. In line 4, the reasoning string is initialized with a template and iteratively updated to generate new attacking prompts. This process repeats for $T$ iterations.